[Ncep.nhc.nco_contacts] TSB Morning Rounds - Thu Oct 20, 2016

TSB Admin nhc.tsbadmin at noaa.gov
Thu Oct 20 14:20:37 UTC 2016


Here is today's summary of NHC computer operations:
--- mello

----------

TAFB:

TAFB:
Update on cumulus performance issues. Mark continues to work with NCO to  
find the proper video driver and card. The problem might persist for a day  
or two.


TSB will test a fix to running clipping scripts on Crappie today.

The experimental rwps data in GFE is not updating. The problem is not local.
HSU:
HSU:
The problems requesting ATCF guidance are no longer an issue.
Other items:

NHC RFCs:


-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20161020/3488c224/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list